Subshilava:

5.5 m 3 m 17.2 m What is the surface area of the prism? 283.8m ^ 2 292.4m ^ 2 325.4m ^ 2 407m ^ 2 What is the lateral surface area of the prism? (1 point) 283.8m ^ 2 • 292.4m ^ 2 325.4m ^ 2 407m ^ 2 Pic below

Subshilava:

1 attachment
Extrinix:

Prisms use the following equation for surface area. \(SA = 2B+ph\) B being the bottom area of the prism, p being the perimeter of the base, and h being the height of the prism. In your case, it'd be: \(SA = 2(17.2\times 3)+((17.2 \times 2)+(3 \times 2))(5.5)\) Solve that and you'll have the Surface Area of your Prism.

Extrinix:

As for the lateral prism, the equation is a bit shorter. \(LSA = Ph\) Yours would be: \(LSA = 2(17.2 + 3) \times 5.5\) Solve that and you will have your Lateral Surface Area of the Prism.